OpenStudy (czesc):

instant fan/medal Which of the following polar equations produces the same graph as the Cartesian equation x^2 - y^2 = 16?

OpenStudy (czesc):

r cos θ = 16 r^2 cos θ = 16 r = 4 r^2 cos 2θ = 16

OpenStudy (dumbcow):

Use following relationships \[x = r \cos \theta\] \[y = r \sin \theta\] \[\cos (2 \theta) = \cos^2 \theta - \sin^2 \theta\]

OpenStudy (czesc):

@dumbcow thank you so much, therefore it would be r^2 cos 2θ = 16 correct? I had never learned those relationships before, thank you again.